A 0.5 hp, 6-pole induction motor is excited by a 3 –phase, 60 Hz source. If the full-load speed is 

1140 rpm, calculate the slip.


1
Expert's answer
2021-09-18T05:06:52-0400

"Ns=\frac{120 \times f}{P} \\

 = \frac{120 \times 60}{6}=1200 \\

slip=\frac{N_s-N_r}{N_s} \\

 =\frac{1200-1140}{1200} \\

 =0.05"